40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 24 01:21:11 2011, kiwamu, Update, SUS, damping test 
    Reply  Mon Sep 26 14:21:07 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps ITMXspectra.pngITMX-sensors.png
       Reply  Mon Sep 26 15:54:46 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps 
          Reply  Mon Sep 26 16:42:08 2011, kiwamu, Update, SUS, ITMX ULSEN : fixed connection.pngITMXsensors.png
             Reply  Wed Sep 28 07:46:42 2011, steve, Update, SUS, Dsub connections at the back of 1X5 are fixed 
Message ID: 5547     Entry time: Mon Sep 26 16:42:08 2011     In reply to: 5546     Reply to this: 5563
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ITMX ULSEN : fixed 

The issue on the ITMX UL sensor has been fixed. It was because of a loose connection in the sensor signal path.

After the fix, the sensor responses completely changed and the suspension became unable to be damped with the new matrix.

At the moment the ITMX suspension is damped by the default input matrix.

we should do the free swinging test once again.

 


(details)

 The loose connection was found on the rear side of the 1X5 rack.

There is an adapter card on the rear side, where the driver and sensor signals are combined into a single cable.

I pushed the sensor cable (bottom right in the picture) and the noise disappeared.

connection.png

Note that I changed the labels on the adapter cards from the old X/Y convention to the new one.

After fixing the loose cable the ITMX suspension became unable to be damped.

So I put the input matrix back to the default and it immediately started damping happily. It means our new matrix is not valid any more.

 

 Here is the latest noise spectra of the ITMX sensors damped with the default input matrix.

As usual all of them are limited by the ADC noise above 20 Hz. (ADC noise is plotted in purple curve)

ITMXsensors.png

 

During the work I also pushed not only ITMX ones but also the cable for the rest of the optics in the adapter cards.

Then PRM became unable to be damped, so it implies the PRM suspension also used to be the same situation.

The input matrix of PRM has been also back to the default.

 

Quote from #5546

Currently the damping of the ITMX suspension is intentionally disabled for the noise investigation.

 

ELOG V3.1.3-